what is the general form of the equation of a circle below
(x-4)^2+y^2=16
1)x^2+y^2-8x-16=16
2)x^2+y^2-8x+16=16
3)x^2+y^2-8x=0
4)x^2+y^2+8x=0

Answer by ewatrrr(24785)

 
Hi,
what is the general form of the equation of a circle below
(x-4)^2+y^2=16
1)x^2+y^2-8x-16=16
2)x^2+y^2-8x+16=16
3)x^2+y^2-8x=0 || (x-4)^2 - 16 + y^2 = 0 or (x-4)^2 + y^2 = 16
4)x^2+y^2+8x=0